OpenStudy (anonymous):

If (4x+2y)^2=64 and xy=2 , what is the value of 4x^2+y^2?

OpenStudy (mertsj):

8

OpenStudy (ash2326):

\[(4x+2y)^2=64\] \[xy=2\] from the first equation, taking 2 out of the bracket, we get \[4(2x+y)^2=64\] dividing both sides by 4 \[(2x+y)^2=16\] now expanding the bracket, we get \[4x^2+y^2+4xy=16\] xy=2 so \[4x^2+y^2+8=16\] subtracting both sides by 8, we get \[4x^2+y^2=8\]